# FVQ: A Large-Scale Dataset and A LMM-based Method for Face Video Quality Assessment

### v1

@@ 5.3. Ablation Study
Effectiveness of Multi-dimensional Features. Finally, we demonstrate the effectiveness of the proposed face-specific features (i.e., portrait features and face embeddings) and temporal features. Note that spatial features are necessary for the quality assessment task, and thus are included in all of our experiments. The last four rows of Table 5 show that the performance of our model improves with the addition of each feature, which highlights the importance of introducing temporal features and face-specific features to the face video quality assessment task.
@@ 6. Conclusion
In this paper, we explore the in-the-wild face video quality assessment problem for the first time. Specifically, we collect the first large-scale FVQA dataset, FVQ-20K, which contains 20,000 face videos with MOS annotations. Along with the FVQ-20K dataset, we propose the first LMM-based FVQA method, FVQ-Rater, which takes advantage of multi-dimensional features including spatial features, temporal features, portrait features, and achieves quality-related rating and scoring through the LoRA-based instruction tuning technique. Extensive experiments on both FVQ-20K and CFVQA datasets demonstrate the superiority of our FVQ-Rater method. We hope that the proposed FVQ-20K dataset, FVQA benchmarks, and FVQ-Rater method will promote in-depth research works on face video quality assessment.

### v2

@@ 5.3. Ablation Study
Effectiveness of Multi-dimensional Features. Finally, we demonstrate the effectiveness of the proposed face-specific features and temporal features. Note that spatial features are necessary for the quality assessment task, and thus are included in all of our experiments. The last four rows of Table 5 show that the performance of our model improves with the addition of each feature, which highlights the importance of introducing temporal features and face-specific features to the FVQA task.
@@ 6. Conclusion
In this paper, we explore the in-the-wild FVQA problem for the first time. Specifically, we collect the first large-scale FVQA dataset, FVQ-20K, which contains 20,000 face videos with MOS annotations. Along with the FVQ-20K dataset, we propose the first LMM-based FVQA method, FVQ-Rater, which takes advantage of multi-dimensional features including spatial features, temporal features, portrait features, and achieves quality-related rating and scoring through the LoRA-based instruction tuning technique. Extensive experiments on both FVQ-20K and CFVQA datasets demonstrate the superiority of our FVQ-Rater method. We hope that the proposed FVQ-20K dataset, FVQA benchmarks, and FVQ-Rater method will promote in-depth research works on face video quality assessment.
Acknowledgements.
This work was supported in part by the National Natural Science Foundation of China under Grants 62401365, 62225112, 62132006, U24A20220, and in part by the China Postdoctoral Science Foundation under Grant Number BX20250411, 2025M773473.